OpenStudy (anonymous):

find the critical numbers of y=x/(x^2+16)

Parth (parthkohli):

Find the derivative and equate it to 0.

OpenStudy (anonymous):

Not sure how to find the derivative

OpenStudy (anonymous):

Could you please explain how I would solve for the Derivative?

OpenStudy (anonymous):

are you in algebra or calculus class? @sharrisf

OpenStudy (anonymous):

calculus @mark_o.

OpenStudy (anonymous):

Differntial Calculus.... u have to use quotient Rule....

OpenStudy (anonymous):

ah ok so you know how to find slope of the tangent? you can do it here \[\lim (x \rightarrow0) \frac{ \Delta y }{ \Delta x }=\lim ( x \rightarrow0) \frac{ f(x+h)-f(x) }{ h } \]

OpenStudy (anonymous):

???? let me help you,what do you think? i need your out put..lol

OpenStudy (anonymous):

Let's put it this way, ever since we started with limits I've been hearing greek and seeing squiggles

OpenStudy (anonymous):

i understand that we all started there..lol

OpenStudy (anonymous):

\[y+\Delta y=f(x+\Delta x)\] then \[\Delta y=f(x+\Delta x) -y\]

OpenStudy (anonymous):

if y=16x^2 then \[y+\Delta y=16(x+\Delta x)^{2}\]

OpenStudy (anonymous):

therefore \[\Delta y=16(x+\Delta x)^{2}-y\] , but y=16x^2 so subs it to the eq

OpenStudy (anonymous):

its now \[\Delta y=16(x+\Delta x)^{2}-16x ^{2}\]

OpenStudy (anonymous):

you can now expand that to \[\Delta y=16(x ^{2}+2x \Delta x +\Delta x ^{2})-16x ^{2}\] expand that more

OpenStudy (anonymous):

\[\Delta y=16x ^{2}+32x \Delta x+16\Delta x ^{2}-16x ^{2}\] do addition or subtraction what ever needed

OpenStudy (anonymous):

so \[\Delta y=32x \Delta x+16\Delta x ^{2}\] thats your delta y or your changed in y

OpenStudy (anonymous):

also \[\frac{ \Delta y }{ \Delta x }=\frac{ 32x \Delta x +16\Delta x ^{2} }{ \Delta x }\] you can divide them

OpenStudy (anonymous):

therefore \[\frac{ \Delta y }{ \Delta x }=\frac{ 32x +16\Delta x }{ \Delta x }\]

OpenStudy (anonymous):

now apply the limit here as x approach to zero \[ \lim_{\Delta x \rightarrow 0}\frac{ \Delta y }{ \Delta x }=32x\]